B) Supposed the probability that a student will pass an Economics exam is 1/2, the probability that he will pass an Economics and Management exam is 3/5, and the probability that he will pass an

Economics or a Management exam is 

4/5. Find the probability that the student will pass a management exam.


1
Expert's answer
2021-07-07T15:12:00-0400
"P(E\cup M)=P(E)+P(M)-P(E\cap M)"

"P(M)=\dfrac{4}{5}+\dfrac{3}{5}-\dfrac{1}{2}=\dfrac{9}{10}"